# Periodic Asynchrony: An On-Policy Approach for Accelerating LLM Reinforcement Learning

### v4

@@ Abstract
Since the introduction of the GRPO algorithm, reinforcement learning (RL) has attracted increasing attention, with growing efforts to reproduce and apply it. However, training efficiency remains a critical challenge. In mainstream RL frameworks, inference and training are typically deployed on the same devices. While this approach reduces costs through resource consolidation, its synchronous execution imposes a computational coupling that prevents concurrent inference and training. In this study, we are returning to the strategy of separating inference and training deployment, and by introducing improvements in the data loader, we transform the conventional synchronous architecture into a periodically asynchronous framework, which allows for demand-driven, independent, and elastic scaling of each component, while the accuracy of the algorithm remains completely equivalent to the synchronization method, with both belonging to the on-policy strategy. It is worth emphasizing that we apply a unified tri-model architecture in the training phase, and we also proposed a shared-prompt attention mask to reduce repetitive computation. In practice, our approach consistently delivers significant end-to-end training efficiency improvements on NPU platforms, indicating its potential for widespread application.
@@ 1 Introduction
Motivated by these limitations, we revisit the training–inference separation strategy and introduce a synchronous framework enhanced with new system-level optimizations. Our design uses a Megatron-style 3D parallel backbone and adds a background producer process that continuously dispatches prompts to inference workers, while the main consumer process retrieves outputs and performs training. Samples from each group are split into micro-batches and trained sequentially, and model weights are synchronized only after the full batch is consumed. This periodic asynchrony achieves purely on-policy asynchronous acceleration and is compatible with any on-policy RL algorithm. The decoupled architecture enables flexible scaling of inference workers, avoiding inference bottlenecks, and for GRPO, we further introduce a shared-prompt attention mask to eliminate redundant computation. Experiments show that when training a 7B-scale model on 16 NPUs, our proposed optimizations achieve over a twofold improvement in end-to-end RL training throughput.
@@ 2 Related Work
For fully asynchronous designs that incorporate off-policy updates, this design effectively trades a small portion of data for off-policy training to prevent transient blocking between inference and learning. Although their experiments demonstrate the effectiveness of this biased setup, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
In contrast, our method takes a complementary algorithmic perspective with several innovations:
  • Algorithm-decoupled periodic asynchrony enabling purely on-policy asynchronous acceleration without sacrificing stability.
  • Decoupling prompt-level inference production from sample-level training consumption, maximizing overlap between the two.
  • A shared-prompt mask for efficient reuse of forward/backward computation on prompt tokens.
  • A distribution-consistent design across policy, old-policy, and reference models, simplifying scaling and system structure.
  • Independent scalability of inference and training instances, preventing bottlenecks in either component.
Together, these designs enable efficient long-context training with substantially fewer computational resources. All components are implemented in project release v5.0.5 (September 5, 2025): https://github.com/janelu9/EasyLLM.
@@ 3 Background
It is well known that when using distributed architectures for large-scale model pre-training or fine-tuning, one typically does not train on an entire batch and immediately update the model parameters with the resulting gradients, as is common in single-GPU training for small models. A full batch usually contains hundreds or even thousands of samples, and directly processing it in large-model training can easily lead to GPU out-of-memory (OOM) errors. Instead, the batch is divided into multiple micro-batches, each containing only one or two samples. These micro-batches are processed sequentially, with a buffer tensor used to accumulate the gradients from each micro-step. The accumulated gradients are applied to update the model parameters only after all micro-batches within the batch have been processed. In practice, these two training strategies are mathematically equivalent.
The gradient update rule for full-batch training is:
$W_{t+1}=W_{t}-\alpha\frac{1}{N}\sum_{n=1}^{N}g_{n},$ (1)
and for micro-batch training, after accumulating gradients over all micro-batches, the update rule is:
$W_{t+1}=W_{t}-\alpha\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}g_{i,j},$ (2)
where $M$ is the total number of micro-batches, $m$ is the number of samples in each micro-batch, and $g_{i,j}$ is the gradient of the $j$-th sample in the $i$-th micro-batch. This demonstrates that both training strategies are mathematically equivalent, but micro-batching offers significant memory savings.
Consider the objective function of GRPO for generating responses from a single prompt:
$J_{\mathrm{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\Big(L_{i}(\pi_{\theta}\|\pi_{\mathrm{old}},A_{i})-\beta D_{KL}^{i}(\pi_{\theta}\|\pi_{\mathrm{ref}})\Big).$ (3)
The PPO-style clipped advantage term can be split into two lines as follows:
@@
$\text{clip}\!\Big(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}|q)},1-\varepsilon,\,1+\varepsilon\Big)A_{i}\Bigg).$ (4)
The KL divergence term is given by:
@@
For a batch of $N$ prompts, the objective becomes:
$J_{\mathrm{batch}}$ $=\frac{1}{N}\sum_{i=1}^{N}J_{\mathrm{GRPO}}^{i}$
$=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{G}\sum_{j=1}^{G}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big)$
$=\frac{1}{NG}\sum_{i=1}^{NG}\Big(L_{i}-\beta D_{KL}^{i}\Big).$ (6)
Define
$M=\frac{NG}{m}.$ (7)
Then the batch objective can be rewritten using micro-batches:
$J_{\mathrm{batch}}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big).$ (8)
This demonstrates that reinforcement learning can also be trained using micro-batches. Compared with pre-training or fine-tuning, a batch now contains $NG$ samples, but GPU memory usage is greatly reduced. Consequently, reinforcement learning can handle sequences as long as those supported during fine-tuning.
@@ 4.1 Unified Tri-Model Architecture
For the third step, mainstream reinforcement learning systems typically require the computation of three types of logits in the forward pass: policy, old policy, and reference. These models share the same architecture, differing only in their weights. In this work, the training phase is implemented using a Megatron-like architecture, where a complete policy model is distributed across multiple computational units according to tensor parallelism and pipeline parallelism scales. During the forward computation of the old policy and reference models, each sub-network managed by the policy model is replicated twice across all computational units—once for computing the old policy logits and once for computing the reference logits. The reference network retains the original model weights, while the old policy network holds weights delayed by one training step relative to the policy model.
Figure 1: Triple Models with Shared Distribution.
The overall architecture of the reinforcement learning system constructed in this work is illustrated in Figure 1. The three types of models share an identical architecture and are distributed in the same parallel layout. This design allows their logits to be computed simultaneously without requiring any additional topology for the reference model. The inference stage, in contrast, is handled by an independent inference engine such as vLLM. Before training each batch, the weights of the policy model are synchronized to all inference engines. The inference engines then receive prompts from the dataset, generate responses, and return them for scoring. Each scored response is concatenated with its corresponding prompt to form a training sample, which is fed into all three models for forward computation. In a single micro-step, we obtain the logits from the policy, old policy, and reference models simultaneously. After all samples in a batch complete their forward passes—and before the policy model’s weights are updated—its current weights are moved into the old policy model’s state_dict.
4.2 Shared-Prompt Attention
In reinforcement learning training, each group of samples shares the same prompt. In fact, during micro-batch training, if the micro-batch size is greater than 1, the prompt can be reused. When the prompt contains many tokens while the generated responses are relatively short, reusing the prompt can significantly reduce redundant computation and GPU memory usage, thereby improving overall training efficiency.
Compared with standard training, when adopting the shared-prompt approach, the following modifications are made:
  1. 1.
    In this formulation, $x_{p}$ denotes the token IDs of the shared prompt, and $x_{r_{1}},x_{r_{2}}$ correspond to the tokens of the individual responses.
  2. 2.
    $p=($ $0,1,\dots,|x_{p}|-1,\ |x_{p}|,\dots,|x_{p}|+|x_{r_{1}}|-1,$
    $|x_{p}|,\dots,|x_{p}|+|x_{r_{2}}|-1)$ (10)
    Here, the notation $|\cdot|$ indicates sequence length.
  3. 3.
    Figure 2: Shared-prompt attention mask.
  4. 4.
All other aspects remain the same as in standard training. This grouped micro-batch approach is generally more computationally efficient than simply concatenating micro-batches along the first axis, because the prompt tokens only need to be computed once, and padding is not required for shorter samples within a micro-batch. The advantage becomes especially pronounced when the prompt is long and the responses are relatively short.
@@ 4.3 Periodic Asynchronous Reinforcement Learning
The main process, acting as the consumer, continuously retrieves completed samples from the queue and feeds them to the three models for training. Under this design, the start time of training is determined by the earliest completed sample among all prompts in the batch. Once all prompts in the batch have been processed, the inference service enters an idle state until the training step finishes and the updated policy weights are synchronized back to the rollout workers. Only then does the next training cycle begin. Therefore, we refer to this design as periodic asynchrony.
@@
In this way, the system constitutes a fully on-policy asynchronous reinforcement learning framework that remains decoupled from the underlying RL algorithm, meaning that no algorithmic modifications are required to ensure correct operation within this asynchronous setup. Algorithm 1 is the pseudocode for the asynchronous reinforcement learning system proposed in this paper.
@@
After that, all samples in the batch are trained sequentially in the order in which they enter the queue. It is clear that when inference speed and training speed are well balanced, the sum of their waiting times is minimized, resulting in optimal end-to-end training throughput on a given hardware configuration. The decoupled inference–training architecture naturally enables such balancing by allowing independent configuration of inference and training instances.
Therefore, in theory, the asynchronous system can achieve nearly twice the end-to-end training speed of the synchronous system.
@@ 5 Implementation
Our implementation is built on the PyTorch training framework, using a 3D-parallel distributed architecture. We incorporate parts of Megatron-Core and DeepSpeed Rasley et al. (2020). On the NPU platform, we additionally use MindSpeed, torch_npu, and optimized operators such as npu_fusion_attention, an accelerated attention kernel supporting custom masks and serving as a counterpart to flash_attention Dao (2023). For inference, we use vLLM, with ascend_vLLM on NPUs. As in fine-tuning, all RL samples are computed at their native sequence lengths for both forward and backward passes without padding, following dynamic-length training.
@@ 6 Experiments
This study conducts comparative experiments on mathematical reasoning tasks, evaluating four reinforcement learning frameworks: MindSpeed-RL, the official NPU training framework, which uses a Megatron backend and adopts a shared-accelerator design for training and inference; VERL, a currently mainstream reinforcement learning framework, which uses an FSDP backend and supports asynchronous execution; the synchronous reinforcement learning framework implemented in this work under a decoupled training–inference design; and the proposed asynchronous reinforcement learning framework. The main goal is to experimentally compare periodic asynchronous and synchronous training, assess the performance of the proposed system against mainstream synchronous frameworks, and evaluate the scalability of the periodic asynchronous system. The primary evaluation metric is end-to-end training throughput, measured by tokens trained per second per device (TPSPD), calculated as the total number of tokens processed divided by the total elapsed time and the number of devices used. Accuracy metrics are provided for reference to verify algorithm correctness, as no modifications were made to the reinforcement learning algorithm.

### v5

@@ Abstract
Since the introduction of the GRPO algorithm, reinforcement learning (RL) has attracted increasing attention for LLM post-training, yet training efficiency remains a critical challenge. In mainstream RL frameworks, inference and training are co-located on the same devices, and their synchronous execution prevents concurrent inference and training. In this work, we revisit the strategy of separating inference and training deployment, and propose a periodically asynchronous framework that transforms synchronous RL training into an asynchronous producer–consumer pipeline. Unlike existing asynchronous approaches that introduce off-policy bias, our design is provably equivalent to its synchronous counterpart, preserving strict on-policy correctness without any algorithmic modifications. We further introduce a unified tri-model architecture and a shared-prompt attention mechanism to support efficient asynchronous execution and reduce redundant computation. Experiments on NPU platforms demonstrate a three- to five-fold improvement in end-to-end training throughput over mainstream RL frameworks, while maintaining fully comparable accuracy, indicating its potential for widespread application.
@@ 1 Introduction
Motivated by these limitations, we revisit the training–inference separation strategy and introduce a synchronous framework enhanced with new system-level optimizations. Our design uses a Megatron-style 3D parallel backbone and adds a background producer process that continuously dispatches prompts to inference workers, while the main consumer process retrieves outputs and performs training. Samples from each group are split into micro-batches and trained sequentially, and model weights are synchronized only after the full batch is consumed. This periodic asynchrony achieves purely on-policy asynchronous acceleration and is compatible with any on-policy RL algorithm. The decoupled architecture enables flexible scaling of inference workers, avoiding inference bottlenecks, and for GRPO, we further introduce a shared-prompt attention mask to eliminate redundant computation. Experiments on NPU platforms show that our proposed framework achieves a three- to five-fold improvement in end-to-end training throughput over mainstream RL frameworks, with near-linear scaling as the number of devices increases.
@@ 2 Related Work
These fully asynchronous designs effectively trade a degree of on-policy correctness for throughput, preventing transient blocking between inference and learning. Although their experiments demonstrate the effectiveness of this approach, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
In contrast, our method takes a complementary algorithmic perspective: we propose a periodically asynchronous framework that preserves strict on-policy correctness without algorithmic modifications, while introducing a unified tri-model architecture and shared-prompt attention mechanism to further improve system efficiency. Unlike existing asynchronous approaches, our design is provably equivalent to synchronous training, offering theoretical guarantees that generalize across on-policy RL algorithms.
@@ 3 Background
We first establish that micro-batch training is mathematically compatible with on-policy RL objectives, as this forms the theoretical foundation for the gradient accumulation scheme used in our asynchronous framework.
In distributed large-scale training, micro-batching is a standard technique that partitions a full batch into smaller micro-batches, accumulating gradients across micro-steps before each parameter update. This is mathematically equivalent to full-batch training while avoiding out-of-memory errors. We extend this equivalence to RL and show it holds for GRPO.
The GRPO objective for a single prompt with $G$ sampled responses is:
$J_{\mathrm{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\Big(L_{i}(\pi_{\theta}\|\pi_{\mathrm{old}},A_{i})-\beta D_{KL}^{i}(\pi_{\theta}\|\pi_{\mathrm{ref}})\Big),$ (1)
where the PPO-style clipped advantage term is defined as:
@@
$\text{clip}\!\Big(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}|q)},1-\varepsilon,\,1+\varepsilon\Big)A_{i}\Bigg),$ (2)
and the KL divergence penalty is:
@@
For a training batch of $N$ prompts, the batch-level objective aggregates over all $NG$ response samples:
$J_{\mathrm{batch}}=\frac{1}{N}\sum_{i=1}^{N}J_{\mathrm{GRPO}}^{i}=\frac{1}{NG}\sum_{k=1}^{NG}\Big(L_{k}-\beta D_{KL}^{k}\Big).$ (4)
Partitioning the $NG$ samples into $M=\lfloor NG/m\rfloor$ micro-batches of size $m$, the objective decomposes as:
$J_{\mathrm{batch}}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{m}\sum_{j=1}^{m}\Big(L_{i,j}-\beta D_{KL}^{i,j}\Big),$ (5)
which is mathematically equivalent to the full-batch update. This confirms that GRPO training is fully compatible with micro-batching: although each batch now contains $NG$ samples rather than $N$, memory consumption remains bounded by the micro-batch size $m$, enabling reinforcement learning over sequences as long as those supported in supervised fine-tuning. This property is directly exploited in our asynchronous framework, where samples arriving from the inference queue are accumulated as micro-batches and trained sequentially without waiting for the full batch to complete.
@@ 4.1 System Overview
For the third step, mainstream reinforcement learning systems typically require the computation of three types of logits in the forward pass: policy, old policy, and reference. These models share the same architecture, differing only in their weights. How these three models are organized and distributed in our system will be detailed in Section 4.2, as their design is closely tied to the requirements of asynchronous execution.
@@ 4.2.1 Asynchronous Execution Mechanism
The main process, acting as the consumer, continuously retrieves completed samples from the queue and feeds them to the training engine. Under this design, the start time of training is determined by the earliest completed sample among all prompts in the batch. Once all prompts in the batch have been processed, the inference service enters an idle state until the training step finishes and the updated policy weights are synchronized back to the rollout workers. Only then does the next training cycle begin. Therefore, we refer to this design as periodic asynchrony.
As implied by the consumer-side design, each training step requires the simultaneous computation of three types of logits—policy, old policy, and reference—for every consumed sample. To address this, we adopt a unified tri-model architecture in which all three models share an identical parallel layout across computational units, implemented on a Megatron-like backbone distributed according to tensor parallelism and pipeline parallelism scales. Each sub-network is replicated twice—once for the old policy logits and once for the reference logits—where the reference network retains the original model weights and the old policy network holds weights delayed by one training step.
Figure 2: Triple Models with Shared Distribution.
As illustrated in Figure 2, this shared distribution allows the logits of all three models to be computed simultaneously in a single micro-step. After all samples in a batch complete their forward passes, the policy model’s current weights are moved into the old policy model’s state_dict before the next parameter update. This design ensures consistent and well-defined weight synchronization at every update boundary, serving as a structural prerequisite for the correctness of the periodic asynchronous framework. Furthermore, since all three models share the same parallel topology, their computational resources are managed uniformly, eliminating the need for separate resource allocation and scheduling across model types and further simplifying system design.
The complete procedure is summarized in Algorithm 1.
@@
4.2.2 Efficiency Analysis
@@
After that, all samples in the batch are trained sequentially in the order in which they enter the queue. It is clear that when inference speed and training speed are well balanced, the sum of their waiting times is minimized, resulting in optimal end-to-end training throughput on a given hardware configuration. Therefore, in theory, the asynchronous system can achieve nearly twice the end-to-end training speed of the synchronous system. Although in rare extreme cases the training time of a single step may be dominated entirely by either inference or training alone, such imbalance can be mitigated through complementary optimizations—for instance, the shared-prompt attention mechanism introduced in Section 4.3 effectively reduces training computation, and the decoupled inference–training architecture naturally enables load balancing by allowing independent scaling of inference and training instances.
4.2.3 Correctness Analysis
As shown in Figure 3(b), the asynchronous system trains samples in a different order from the synchronous system, determined by inference completion time rather than the original batch order. We now formally establish that neither the on-policy correctness nor the gradient equivalence is compromised by this reordering.
Proposition 1 (Periodic Weight Consistency). All rollout samples within the same batch are generated by the same policy $\pi_{\theta_{t}}$, i.e.:
$\forall\,i\in\{1,\ldots,B\}:\quad o_{i}\sim\pi_{\theta_{t}}(\cdot\mid p_{i}).$ (8)
Proof. Line 3 synchronizes all inference workers to $\theta_{t}$ before any rollout begins, and no weight update is performed until Line 21, after all $B$ samples have been consumed. Therefore, every $o_{i}$ is sampled exclusively from $\pi_{\theta_{t}}$, satisfying the on-policy condition identically to a synchronous system.
Having established on-policy sample consistency, we next show that the order in which these samples are consumed during training does not affect the gradient update.
Proposition 2 (Gradient Permutation Invariance). Let $\sigma$ be any permutation of the $NG$ training samples across micro-batches. The accumulated gradient is invariant to $\sigma$:
$\nabla_{\theta}J_{\mathrm{batch}}=\frac{1}{NG}\sum_{k=1}^{NG}\nabla_{\theta}\Big(L_{k}-\beta D_{KL}^{k}\Big).$ (9)
Proof. The double summation $\sum_{i=1}^{M}\sum_{j=1}^{m}$ enumerates all $NG$ samples regardless of their ordering. The result follows directly from the commutativity of finite summation.
This guarantees that the periodic asynchronous system is algorithmically equivalent to its synchronous counterpart, with any on-policy RL algorithm remaining correct under asynchronous execution without modification. Combining Proposition 1 and Proposition 2, we arrive at the main theorem of this section.
Theorem 1 (Equivalence of Asynchronous and Synchronous Updates). At each training iteration $t$, the periodic asynchronous system produces an identical parameter update to its synchronous counterpart:
$\Delta\theta_{\mathrm{async}}^{t}=\Delta\theta_{\mathrm{sync}}^{t}.$ (10)
Consequently, the two systems follow exactly the same training trajectory $\{\theta_{t}\}$, confirming that periodic asynchrony is purely on-policy and algorithmically equivalent to synchronous RL training.
We can thus conclude that the proposed periodic asynchronous framework constitutes a fully on-policy asynchronous reinforcement learning system that is algorithmically decoupled from the underlying RL algorithm, requiring no modifications to ensure correct operation. Its convergence and stability are identical to those of any synchronous on-policy method it is built upon.
4.3 Shared-Prompt Attention
In GRPO-based reinforcement learning, all samples within a group are generated from the same prompt, making it possible to share the prompt computation across responses within a micro-batch, which significantly reduces redundant computation and memory consumption when prompts are long relative to responses.
Compared with standard training, the shared-prompt approach introduces four targeted modifications to the input construction, positional encoding, attention mask, and loss computation respectively:
  1. 1.
    where $x_{p}$ denotes the token IDs of the shared prompt, and $x_{r_{1}},x_{r_{2}}$ correspond to the tokens of the individual responses.
  2. 2.
    $p=($ $0,1,\dots,|x_{p}|-1,\ |x_{p}|,\dots,|x_{p}|+|x_{r_{1}}|-1,$
    $|x_{p}|,\dots,|x_{p}|+|x_{r_{2}}|-1)$ (12)
    where $|\cdot|$ denotes sequence length.
  3. 3.
    Figure 4: Shared-prompt attention mask.
  4. 4.
Correctness Analysis.
By construction of the shared-prompt mask (Step 3), each response token attends only to the prompt and its own preceding tokens, with cross-response attention fully masked. Therefore, the hidden states and gradients of each response are determined solely by $x_{p}$ and $x_{r_{k}}$, giving:
$\nabla_{\theta}\mathcal{L}_{\text{shared}}=\sum_{k=1}^{K}\nabla_{\theta}\mathcal{L}_{k},$ (14)
where $\mathcal{L}_{k}$ denotes the loss on the $k$-th response under standard training. This confirms that the shared-prompt approach is mathematically equivalent to standard per-sample training, introducing no approximation or bias into the training objective.
Complexity Analysis.
Let $L_{p}$, $L_{r}$, and $K$ denote the prompt length, average response length, and number of responses per micro-batch, respectively. In standard training, the prompt tokens are redundantly computed $K$ times, giving a total attention complexity of $\mathcal{O}(K(L_{p}+L_{r})^{2})$. Under the shared-prompt approach, the attention computation decomposes into three parts according to the mask structure: prompt self-attention $\mathcal{O}(L_{p}^{2})$, response-to-prompt attention $\mathcal{O}(KL_{r}L_{p})$, and response self-attention $\mathcal{O}(KL_{r}^{2})$, giving a total of $\mathcal{O}(L_{p}^{2}+KL_{r}(L_{p}+L_{r}))$. The reduction ratio between the two approaches is:
$\rho=\frac{L_{p}^{2}+KL_{r}(L_{p}+L_{r})}{K(L_{p}+L_{r})^{2}}$ (15)
When $L_{p}\gg L_{r}$, $\rho$ approaches $\frac{1}{K}$, yielding an approximately $K$-fold reduction in attention computation. Additionally, no padding is required for shorter responses within a micro-batch, further improving hardware utilization. These gains are therefore most pronounced in long-prompt, short-response settings, which are common in complex reasoning tasks with GRPO.
@@ 5 Implementation
Our implementation is built on the PyTorch training framework, using a 3D-parallel distributed architecture. We incorporate parts of Megatron-Core and DeepSpeed Rasley et al. (2020). On the NPU platform, we additionally use MindSpeed, torch_npu, and optimized operators such as npu_fusion_attention, an accelerated attention kernel supporting custom masks and serving as a counterpart to flash_attention Dao (2023). For inference, we use vLLM, with ascend_vLLM on NPUs.
As in fine-tuning, all RL samples are computed at their native sequence lengths for both forward and backward passes without padding, following dynamic-length training. The inference-to-training instance ratio is configurable and set to 1:4 in all experiments, chosen to balance inference and training throughput on the Ascend-910B platform.
All components described above are publicly available in project release v5.0.5 (September 5, 2025) at https://github.com/janelu9/EasyLLM.
@@ 6 Experiments
We evaluate four reinforcement learning frameworks on mathematical reasoning tasks: MindSpeed-RL, an official NPU training framework using a Megatron backend with shared-accelerator design; VERL, a mainstream framework using an FSDP backend with asynchronous execution; our synchronous baseline under a decoupled training–inference design; and our proposed asynchronous framework. The primary evaluation metric is end-to-end training throughput, measured by tokens trained per second per device (TPSPD). Accuracy metrics are provided as reference to verify that no degradation is introduced by our algorithmic design.
6.1 Model and Environment Configuration
The models compared in this section are Qwen2.5-7B-Instruct Team and others (2024), Qwen3-8B Yang et al. (2025), and DeepSeek-R1-Distill-Qwen-32B DeepSeek-AI (2025). The training data comes from the publicly available math problem datasets DeepScaleR Luo and others (2025) and GSM8K Cobbe et al. (2021). All reinforcement learning experiments use the GRPO algorithm. The accuracy test set is AIME24 Zhang and Math-AI (2024).